# Ophiuchus: Scalable Modeling of Protein Structures through Hierarchical Coarse-graining SO(3)-Equivariant Autoencoders

### v1

@@ 5 Conclusion and Future Work
Appendix A All-Atom Decoding
Given a latent representation at the residue-level, $(\mathbf{X},\mathbf{V}^{0:{l_{\max}}})_{i}$, we take $\mathbf{X}_{i}$ directly as $\hat{\mathbf{X}}_{i}^{\alpha}$ to obtain the $\mathbf{C}_{\alpha}$ position, and transform hidden scalar representations $\mathbf{V}_{i}^{l=0}$ into categorical logits over $\mathcal{R}$ to decode $\hat{\mathbf{R}}_{i}$. We then transform $\mathbf{V}_{i}^{l>0}$ to obtain the relative position vectors $\hat{\mathbf{X}}^{\ast}_{i}$. The orderable atoms are obtained from $\mathbf{V}^{l=1}$ while the unorderable pairs are obtained from $(\mathbf{V}^{l=1},\mathbf{V}^{l=2})$. To generate an unorderable pair, we treat $\mathbf{V}^{l=1}$ as the center position of the symmetric atoms – that is, a "mean vector". We then compute the square root of $\mathbf{V}^{l=2}$ to establish the relative difference between the two vectors in the unorderable pair. This method proves effective because the sign of the square root is inherently ambiguous. Both positive and negative square root values yield the same $\mathbf{V}^{l=2}$, aligning perfectly with our requirement for the vectors to be unorderable.
Appendix B Details on Self-Interaction
@@
The objective of our Self-Interaction module is to function exclusively based on $\mathbf{V}^{0:{l_{\max}}}$, while concurrently mixing irreducible representations across various $l$ values. To accomplish this, we calculate the tensor product of the representation with itself; this operation is termed the "tensor square" and is symbolized by $\left(\mathbf{V}^{0:{l_{\max}}}\right)^{\otimes 2}$. As the channel dimensions expand, the computational complexity tied to the tensor square increases quadratically. In the case of latent diffusion models, we employ the tensor square in its original, unmodified form, accepting the computational cost as a trade-off for maintaining full expressivity. Contrastingly, for autoencoder models, we reduce the computational load by dividing the representation into smaller, independent segments along the channel dimension prior to executing the tensor square. The segmented results are subsequently concatenated to generate the updated representation $\mathbf{V}^{0:{l_{\max}}}$.
Appendix C Transpose Sequence Convolution
@@
Given a single, coarse anchor and representation $\{\mathbf{P},\mathbf{V}^{l}\}$, we first map $\{V^{l}\}$ into $K$ new representations linearly, reshaping it appropriately. We then similarly map $\mathbf{V}^{l=1}$ linearly into relative positioning vectors $[\Delta\mathbf{P}]^{K}$, anchoring them in the previous coordinate to obtain $K$ new coordinates, $[\mathbf{P}]^{K}\leftarrow\mathbf{P}+[\Delta\mathbf{P}]^{K}$:
Input: Kernel Size $K$ and Stride $S$.
Input: The Perceptive Field of Varying Size $P$ ($\in\{1,2\}$ for $K=3$ and $S=2$)
Input: Latent Representations $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})_{1:P}$
Input: Relative Indices $\Delta_{1:P}\in[-\frac{K}{2},\frac{K}{2}]$
$\mathbf{V}^{0:{l_{\max}}}_{1:P}\leftarrow\textrm{Norm}\left(\mathbf{V}^{0:{l_{\max}}}_{1:P}\right)$  normal-$\triangleright$ Normalize $l$-Wise
@@
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\frac{1}{P}\sum_{p=1}^{P}\textrm{Linear}\left(\Delta_{p},\mathbf{V}^{0:{l_{\max}}}_{p}\right)$  normal-$\triangleright$ Different Parameters For Each $\Delta$
@@
$\mathbf{P}\leftarrow\frac{1}{P}\sum_{p=1}^{P}(\mathbf{P}_{p}+\textrm{Linear}\left(\Delta_{p},\mathbf{V}^{l=1}_{p}\right))$ return $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$

### v2

@@ 5 Conclusion and Future Work
Appendix A Architecture Details
A.1 All-Atom Representation
A canonical ordering of the atoms of each residue enables the local geometry to be described in a stacked array representation, where each feature channel corresponds to an atom. To directly encode positions, we stack the 3D coordinates of each atom. The coordinates vector behaves as the irreducible-representation of $SO(3)$ of degree $l=1$. The atomic coordinates are taken relative to the $\mathbf{C}_{\alpha}$ of each residue. In practice, for implementing this ordering we follow the atom14 tensor format of SidechainNet [King & Koes (2021)], where a vector $P\in\mathbb{R}^{14\times 3}$ contains the atomic positions per residue. In Ophiuchus, we rearrange this encoding: one of those dimensions, the $\mathbf{C}_{\alpha}$ coordinate, is used as the absolute position; the 13 remaining 3D-vectors are centered at the $\mathbf{C}_{\alpha}$, and used as geometric features. The geometric features of residues with fewer than 14 atoms are zero-padded (Figure 4).
Still, four of the standard residues (Aspartic Acid, Glutamic Acid, Phenylalanine and Tyrosine) have at most two pairs of atoms that are interchangeable, due to the presence $180^{\circ}$-rotation symmetries [Jumper et al. (2021)]. In Figure 5, we show how stacking their relative positions leads to representations that differ even when the same structure occurs across different rotations of a side-chain. To solve this issue, instead of stacking two 3D vectors ($\mathbf{P}^{*}_{u},\mathbf{P}^{*}_{v}$), our method uses a single $l=1$ vector $\mathbf{V}^{l=1}_{\textrm{center}}=\frac{1}{2}(\mathbf{P}^{*}_{v}+\mathbf{P}^{*}_{u})$. The mean makes this feature invariant to the $(v,u)$ atomic permutations, and the resulting vector points to the midpoint between the two atoms. To fully describe the positioning, difference $(\mathbf{P}^{*}_{v}-\mathbf{P}^{*}_{u})$ must be encoded as well. For that, we use a single $l=2$ feature $\mathbf{V}^{l=2}_{\textrm{diff}}=Y_{l=2}\big{(}\frac{1}{2}(\mathbf{P}^{*}_{v}-\mathbf{P}^{*}_{u})\big{)}$. This feature is produced by projecting the difference of positions into a degree $l=2$ spherical harmonics basis. Let $x\in\mathbb{R}^{3}$ denote a 3D vector. Then its projection into a feature of degree $l=2$ is defined as:
$Y_{2}(x)=\left[\sqrt{15}\cdot x_{0}\cdot x_{2},\sqrt{15}\cdot x0\cdot x1,\frac{\sqrt{5}}{2}\left(-x_{0}^{2}+2x_{1}^{2}-x_{2}^{2}\right),\sqrt{15}\cdot x_{1}\cdot x_{2},\frac{\sqrt{15}}{2}\cdot(-x_{0}^{2}+x_{2}^{2})\right]$
Where each dimension of the resulting term is indexed by the order $m\in[-l,l]=[-2,2]$, for degree $l=2$. We note that for $m\in\{-2,-1,1\}$, two components of $x$ directly multiply, while for $m\in\{0,2\}$ only squared terms of $x$ are present. In both cases, the terms are invariant to flipping the sign of $x$, such that $Y_{2}(x)=Y_{2}(-x)$. Equivalently, $\mathbf{V}^{l=2}_{\textrm{diff}}$ is invariant to reordering of the two atoms $(v,u)\rightarrow(u,v)$:
$\mathbf{V}^{l=2}_{\textrm{diff}}=Y_{l=2}\left(\frac{1}{2}(\mathbf{P}^{*}_{v}-\mathbf{P}^{*}_{u})\right)=Y_{l=2}\left(\frac{1}{2}(\mathbf{P}^{*}_{u}-\mathbf{P}^{*}_{v})\right)$
In Figure 5, we compare the geometric latent space of a network that uses this permutation invariant encoding, versus one that uses naive stacking of atomic positions $\mathbf{P}^{*}$. We find that Ophiuchus correctly maps the geometry of the data, while direct stacking leads to representations that do not reflect the underlying symmetries.
Figure 4: Geometric Representations of Protein Residues. (a) We leverage the bases of spherical harmonics to represent signals on SO(3). (b-d) Examples of individual protein residues encoded in spherical harmonics bases: residue label is directly encoded in a order $l=0$ representation as a one-hot vector; atom positions in a canonical ordering are encoded as $l=1$ features; additionally, unorderable atom positions are encoded as $l=1$ and $l=2$ features that are invariant to their permutation flips. In the figures, we displace $l=2$ features for illustrative purposes – in practice, the signal is processed as centered in the $\mathbf{C}_{\alpha}$. (b) Tryptophan is the largest residue we consider, utilizing all dimensions of $n_{\textrm{max}}=13$ atoms in the input $\mathbf{V}^{l=1}_{\textrm{ordered}}$. (c) Tyrosine needs padding for $\mathbf{V}^{l=1}$, but produces two $\mathbf{V}^{l=2}$ from its two pairs of permutable atoms. (d) Aspartic Acid has a single pair of permutable atoms, and its $\mathbf{V}^{l=2}$ is padded.
Figure 5: Permutation Invariance of Side-chain Atoms in Stacked Geometric Representations. (a) We provide an example in which we rotate the side-chain of Tyrosine by $2\pi$ radians while keeping the ordering of atoms fixed. Note that the structure is the same after rotation by $\pi$ radians. (b) An SO(3)-equivariant network may stack atomic relative positions in a canonical order. However, because of the permutation symmetries, the naive stacked representation will lead to latent representations that are different even when the data is geometrically the same. To demonstrate this, we plot two components ($m=-1$ and $m=0$) of an internal $\mathbf{V}^{l=1}$ feature, while rotating the side-chain positions by $2\pi$ radians. This network represents the structures rotated by $\theta=0$ (red) and $\theta=\pi$ (cyan) differently despite having exactly the same geometric features. (c-d) Plots of internal $\mathbf{V}^{l=1,2}$ of Ophiuchus which encodes the position of permutable pairs jointly as an $l=1$ center of symmetry and an $l=2$ difference of positions, resulting in the same representation for structures rotated by $\theta=0$ (red) and $\theta=\pi$ (cyan).
A.2 All-Atom Decoding
Given a latent representation at the residue-level, $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$, we take $\mathbf{P}$ directly as the $\mathbf{C}_{\alpha}$ position of the residue. The hidden scalar representations $\mathbf{V}^{l=0}$ are transformed into categorical logits $\bm{\ell}\in\mathbb{R}^{|\mathcal{R}|}$ to predict the probabilities of residue label $\hat{\mathbf{R}}$. To decode the side-chain atoms, Ophiuchus produces relative positions to $\mathbf{C}_{\alpha}$ for all atoms of each residue. During training, we enforce the residue label to be the ground truth. During inference, we output the residue label corresponding to the largest logit value.
Relative coordinates are produced directly from geometric features. We linearly project $\mathbf{V}^{l=1}$ to obtain the relative position vectors of orderable atoms $\hat{\mathbf{V}}^{l=1}_{\textrm{ordered}}$. To decode positions $(\hat{\mathbf{P}}^{*}_{v},\hat{\mathbf{P}}^{*}_{u})$ for an unorderable pair of atoms $(v,u)$, we linearly project $\mathbf{V}^{l>0}$ to predict $\hat{\mathbf{V}}^{l=1}_{\textrm{center}}$ and $\hat{\mathbf{V}}^{l=2}_{\textrm{diff}}$. To produce two relative positions out of $\hat{\mathbf{V}}^{l=2}_{\textrm{diff}}$, we determine the rotation axis around which the feature rotates the least by taking the left-eigenvector with smallest eigenvalue of $\mathcal{X}^{l=2}\hat{\mathbf{V}}^{l=2}_{\textrm{diff}}$, where $\mathcal{X}^{l=2}$ is the generator of the irreducible representations of degree $l=2$. We illustrate this procedure in Figure 6 and explain the process in detail in the caption. This method proves effective because the output direction of this axis is inherently ambiguous, aligning perfectly with our requirement for the vectors to be unorderable.
Figure 6: Decoding Permutable Atoms. Sketch on how to decode a double-sided arrow ($\mathbf{V}^{l=2}$ signals) into two unordered vectors ($\hat{\mathbf{P}}^{*}_{u}$, $\hat{\mathbf{P}}^{*}_{v}$). $Y^{l=2}:\mathbb{R}^{3}\to\mathbb{R}^{5}$ can be viewed as a 3 dimensional manifold embed in a 5 dimensional space. We exploit the fact that the points on that manifold are unaffected by a rotation around their pre-image vector. To extend the definition to all the points of $\mathbb{R}^{5}$ (i.e. also outside of the manifold), we look for the axis of rotation with the smallest impact on the 5d entry. For that, we compute the left-eigenvector with the smallest eigenvalue of the action of the generators on the input point: $\mathcal{X}^{l=2}\hat{\mathbf{V}}^{l=2}_{\textrm{diff}}$. The resulting axis is used as a relative position ±$\pm\Delta\mathbf{P}^{*}_{v,u}$ between the two atoms, and is used to recover atomic positions through $\mathbf{P}^{*}_{v}=\mathbf{V}^{l=1}_{\textrm{center}}+\Delta\mathbf{P}^{*}_{v,u}$ and $\mathbf{P}^{*}_{u}=\mathbf{V}^{l=1}_{\textrm{center}}-\Delta\mathbf{P}^{*}_{v,u}$.
Figure 7: Building Blocks of Self-Interaction. (a) Self-Interaction updates only SO(3)-equivariant features, which are represented as a $D$ channels each with vectors of degree $l$ up to degree $l_{\textrm{max}}$. (b) A roto-translational equivariant linear layer transforms only within the same order $l$. (c-d) we use the tensor square operation to interact features across different degrees $l$. We employ two instantiations of this operation. (c) The Self-Interaction in autoencoder models applies the square operation within the same channel of the representation, interacting features with themselves across $l$ of same channel dimension $d\in[0,D]$. (d) The Self-Interaction in diffusion models chunks the representation in groups of channels before the square operation. It is more expressive, but imposes a harder computational cost.
A.3 Details of Self-Interaction
@@
The objective of our Self-Interaction module is to function exclusively based on the geometric features $\mathbf{V}^{0:{l_{\max}}}$, while concurrently mixing irreducible representations across various $l$ values. To accomplish this, we calculate the tensor product of the representation with itself; this operation is termed the "tensor square" and is symbolized by $\left(\mathbf{V}^{0:{l_{\max}}}\right)^{\otimes 2}$. As the channel dimensions expand, the computational complexity tied to the tensor square increases quadratically. To solve this computational load, we instead perform the square operation channel-wise, or by chunks of channels. Figures 7.c and 7.d illustrate these operations. After obtaining the squares from the chunked or individual channels, the segmented results are subsequently concatenated to generate an updated representation $\mathbf{V}^{0:{l_{\max}}}$, which is transformed through a learnable linear layer to the output dimensionality.
A.4 Non-Linearities
To incorporate non-linearities in our geometric representations $\mathbf{V}^{l=0:l_{\textrm{max}}}$, we employ a similar roto-translation equivariant gate mechanism as described in Equiformer [Liao & Smidt (2023)]. This mechanism is present at the last step of Self-Interaction and in the message preparation step of Spatial Convolution (Figure (2)). In both cases, we implement the activation by first isolating the scalar representations $\mathbf{V}^{l=0}$ and transforming them through a standard MultiLayerPerceptron (MLP). We use the SiLu activation function [Elfwing et al. (2017)] after each layer of the MLP. In the output vector, a scalar is produced for and multiplied into each channel of $\mathbf{V}^{l=0:l_{\textrm{max}}}$.
Figure 8: Convolutions of Ophiuchus. (a) In the Spatial Convolution, we update the feature representation $\mathbf{V}_{i}^{0:l_{\textrm{max}}}$ and position $\mathbf{P}_{i}$ of a node $i$ by first aggregating messages from its $k$ nearest-neighbors. The message is composed out of the neighbor features $\mathbf{\tilde{V}}_{1:k}^{0:l_{\textrm{max}}}$ and the relative position between the nodes $\mathbf{P}_{i}-\tilde{\mathbf{P}}_{1:k}$. After updating the features $\mathbf{V}_{i}^{0:l_{\textrm{max}}}$, we project vectors $\mathbf{V}_{i}^{l=1}$ to predict an update to the position $\mathbf{P}_{i}$. (b) In the Sequence Convolution, we concatenate the feature representations of sequence neighbors $\mathbf{\tilde{V}}_{1:K}^{0:l_{\textrm{max}}}$ along with spherical harmonic signals of their relative positions $\mathbf{P}_{i}-\mathbf{P}_{j}$, for $i\in[1,K]$, $j\in[1,K]$. The concatenated feature vector is then linearly projected to the output dimensionality. To produce a coarsened position, each $\mathbf{\tilde{V}}_{1:K}^{0:l_{\textrm{max}}}$ produces a score that is used as weight in summing the original positions $\mathbf{P}_{1:K}$. The Softmax function is necessary to ensure the sum of the weights is normalized.
A.5 Roto-Translation Equivariance of Sequence Convolution
A Sequence Convolution kernel takes in $K$ coordinates $\mathbf{P}_{1:K}$ to produce a single coordinate $\mathbf{P}=\sum_{i=1}^{K}w_{i}\mathbf{P}_{i}$. We show that these weights need to be normalized in order for translation equivariance to be satisfied. Let $T$ denote a 3D translation vector, then translation equivariance requires:
$(\mathbf{P}+T)=\sum_{i=1}^{K}w_{i}(\mathbf{P}_{i}+T)=\sum_{i=1}^{K}w_{i}\mathbf{P}_{i}+\sum_{i=1}^{K}w_{i}T\;\;\rightarrow\;\;\sum_{i=1}^{K}w_{i}=1$
Rotation equivariance is immediately satisfied since the sum of 3D vectors is a rotation equivariant operation. Let $R$ denote a rotation matrix. Then,
$(R\mathbf{P})=R\sum_{i=1}^{K}w_{i}(\mathbf{P}_{i})=\sum_{i=1}^{K}w_{i}(R\mathbf{P}_{i})$
A.6 Transpose Sequence Convolution
@@
Given a single coarse anchor position and features representation $(\mathbf{P},\mathbf{V}^{l=0:l_{\textrm{max}}})$, we first map $\mathbf{V}^{l=0:l_{\textrm{max}}}$ into a new representation, reshaping it by chunking $K$ features. We then project $\mathbf{V}^{l=1}$ and produce $K$ relative position vectors $\Delta\mathbf{P}_{1:K}$, which are summed with the original position $\mathbf{P}$ to produce $K$ new coordinates.
Input: Kernel Size $K$
Input: Latent Representations $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$
@@
$\Delta\mathbf{P}_{1:K}\leftarrow\textrm{Linear}\left(\mathbf{V}^{l=1}\right)$   normal-$\triangleright$ Predict $K$ relative positions
@@
$\mathbf{P}_{1:K}\leftarrow\mathbf{P}+\Delta\mathbf{P}_{1:K}$ $\mathbf{V}^{l=0:l_{\textrm{max}}}_{1:K}\leftarrow\textrm{Reshape}_{K}\left(\textrm{Linear}(\mathbf{V}^{l=0:l_{\textrm{max}}})\right)$   normal-$\triangleright$ Split the channels in $K$ chunks
return $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})_{1:K}$